Assume TR(Q) = 150 + 500-50° and TC(Q) = 100 + 6Q Create a table in Excel showing TR(Q), MR (0), TC(Q), MC (Q), π (Q), and marginal profit, as O goes from 0 to 10 in increments of 1 b. Identify the level of Q that maximizes profit. c. At the level of Q that maximizes profit, what is the relation between MR and MC? d. At the level of Q that maximizes profit, what is marginal profit? e. Check your answer in part b using calculus
Answer #1

Marginal revenue function = d(TR)/dQ

MR = d(150+50Q-5Q2)/dQ = 50-10Q

MC = d(TC)/dQ

MC = d(100+6Q)/dQ = 6

Q TR=150+50Q - 5Q2 TC =100+6Q MC MR Marginal Profit= MR - MC
0 150 100 - - -
1 195 106 6 45 39
2 230 112 6 35 29
3 255 118 6 25 19
4 270 124 6 15 9
5 275 130 6 5 -1
6 270 136 6 -5 -11
7 255 142 6 -15 -21
8 230 148 6 -25 -31
9 195 154 6 -35 -41
10 150 160 6 -45 -51
Profit (TR-TC)
50
89
118
137
146
145
134
113
82
41
-10

B) At Q = 4 the profit is maximum

C) At the level of Q = 4 the marginal revenue is decreasing and the marginal cost is constant.

D) At Q = 4 the marginal profit is 9

E) When MR = MC we get the maximum quantity which maximize profit

MR = MC

50-10Q = 6

44 = 10Q

Q = 4.4

So the value is approximately equal to 4